OpenStudy (anonymous):

A 50ml of10% acid solution want to make 25% acid solution. How much pure acid do i add?

OpenStudy (anonymous):

\[.1\times 50+x=.25(50+x)\] solve for x

OpenStudy (anonymous):

\[5+x=12.5+.25x\] \[.75x=7.5\] \[x=7.5\div .75=10\]